Amar Monroe NY on a smooth floor as initial velocity of 0.4 metre per second is the floor office and retardation of 0.02 metre per second square calculate the time it will take to come rest

Grade:9

1 Answers

hero
6 years ago
v=u+at
v= 0 , u=0.4 , a=-0.02
using these
0=0.4 +(-0.02)×t
t=0.4/0.02
t= 20 seconds 
hope it helps
for any doubt plz feel free ask again.